A rectangular dog pen which is 5 times as long as it is wide
Answer by ewatrrr(24785)

 
Hi,

A rectangular dog pen which is 5 times as long as it is wide
P = 2L + 2w = 2(5w) + 2w = 12w
A = Lw = (5w)(w)= 5w^2